OpenStudy (_haii_):

h e l p MEDAL + FAN Match the choices to the name of the form of each linear equation. a) 𝑦 − 15 = 8/9 (9𝑥 − 4) b) 𝑦 − 12 = − 3/2 (𝑥 + 1/2 ) c) 𝑦 = 13/14 𝑥 − 1/8 d) 7𝑥 + 20𝑦 = 78 e) 9𝑥 = 2𝑦 − 1 f) 𝑦 + 71 = 7/8 (𝑥 + 3) g) 𝑦 = −2𝑥 + 3 h) 7𝑥 + 1/7 𝑦 = 19 A. Point-Slope Form B. Slope-Intercept Form C. Standard Form D. None

OpenStudy (will.h):

Your equations are not clear. There are weird symbols

OpenStudy (will.h):

But anyway I'll give you examples that will help you in determining your answers Point slope formula: y - y1 = m(x - x1) Slope intercept form: y= mx + b Standard form: y + x = a number

OpenStudy (mathmale):

Will is right, although I'd write "standard form" as Ax + By = C. Look through the given equations and choose the form (i. e., give the name) of the equation of a straight line that is closest.

OpenStudy (mathmale):

By "𝑦 − 15 = 8/9 (9𝑥 − 4) " you seem to mean 𝑦 − 15 = (8/9) (9𝑥 − 4) . You should enclose fractions such as this 8/9 inside parentheses for clarity. Your 𝑦 − 15 = 8/9 (9𝑥 − 4) resembles \[y-y _{1}=m(x-x _{1}). \]

OpenStudy (mathmale):

What's the name of that form?

OpenStudy (mathmale):

Continue with the rest of the problems.